The table below shows the values of selected macroeconomics variables over a period:

YEAR 1 (R-Billion) YEAR 2 (R-Billion)


Investment 200 220




Saving 180 190


Export 100 110


Imports 120 140


Government Expenditure 150 160


Taxation 150 160


Equilibrium National Income 1 800 2 000




(a) Calculate the value of the multiplier for this economy.


(b) Should the full employment level of income be R-B2 255, by how much should government change its spending to reach this level of income during the next year given the multiplier has not changed.


(c) Evaluate whether or not this policy approach is effective in real life in achieving the desired level of GDP.



1
Expert's answer
2021-07-28T07:36:01-0400

(a) value of the multiplier .


Found by dividing equilibrium national income by government spending.


"x=\frac {180 0}{150}"


=12


The value of the multiplier is therefore 12 times



(b)

Level of income=2255

Multiplier =12

"12=\frac{2255}{x}"

"12x=2255"

"x=\frac {2255}{12}"

"x=187.9"

The government should thus increase its spending by 27.9

"187.9-150=27.9"




(c) This policy is effective in raising GDP to the desired level. This is because it directly relates the  equilibrium national level of revenue to government spending.